# Stacky Batyrev–Manin conjecture and modular curves

### v1

@@ 2.3. Sectors of some Deligne–Mumford curves
Consider a vector bundle $V$ on $X$. Given an algebraically closed $\overline{F}$-field $K$ an $x\in(\mathcal{J}_{0}X)(K)$ corresponding to $(x^{\prime},\mu_{n}\hookrightarrow\operatorname{Aut}(x))$, we denote by $\rho(x)$ the induced action of $\mu_{n}$ on $V(x)$. We define
@@
If $V=T_{X}$, we write $\operatorname{age}$ for $\operatorname{age}_{V}$.
@@
Let $k$ be an algebraically closed field $\overline{F}$-field. A finite dimensional $k$-representation of $\widehat{\mu}$ is an $k$-homomorphism $\rho:\widehat{\mu}\to\operatorname{GL}_{r}$, for certain $r\geq 1$ called the dimension of representation, which factorizes through the canonical homomorphism $\widehat{\mu}\to\mu_{\ell}$ for certain $\ell\geq 1$. We let $\rho_{\ell}:\mu_{\ell}\to\operatorname{GL}_{r}$ be the induced representation. We define
@@
Let $k$ be an algebraically closed field $\overline{F}$-field such that there exists a $k$-point $x\in\mathcal{X}$. The point $x$ corresponds to a point $x\in X(k)$ and a $k$-homomorphism $\widehat{\mu}\to\operatorname{Aut}(x).$ We have a commutative diagram
$\operatorname{Aut}(x)$$\operatorname{GL}(f^{*}V(x))=\operatorname{GL}(V(f(x))$$\operatorname{Aut}(f(x))$$\operatorname{GL}(V(f(x))$$\scriptstyle{=}$
@@
Suppose we have $f:X\to Y$ as in Lemma 2.12 which is also assumed étale at a neighborhood of a point $x$ of $X$. Let $\mathcal{X}\in\pi_{0}\mathcal{J}_{0}X$ be any sector above $x\in|X|$, and let $\mathcal{Y}=(\pi_{0}\mathcal{J}_{0}f)(\mathcal{X})$. Then
@@
One has that $f$ is étale, then the tangent bundles satisfy $f^{*}T_{Y}=T_{X}$. The corollary follows from Lemma 2.12. ∎
@@ 2.3.3. Weighted projective stacks
Let $n\geq 1$ be an integer. Suppose that $a_{1},\dots,a_{n}\geq 1$ are given integers. We consider the weighted projective stack $\mathcal{P}(a_{1},\dots,a_{n}):=[(\mathbb{A}^{n}-\{0\})/\mathbb{G}_{m}]$. For any $F$-scheme $S$ and any morphism $f:S\to\mathcal{P}(a_{1},\dots,a_{n})$, we have a canonical identification $\operatorname{Aut}(f)=\mu_{n(f)}$ for certain $n(f)|\operatorname{lcm}(a_{i}).$ We will fix a bijection bijection
@@
Here, for $n|k$, we identify $\theta:\widehat{\mu}\to\mu_{k}$ and $\widehat{\mu}\to\mu_{n}$ which factorizes through $\theta$. If $\widehat{\mu}\to\mu_{n}$ is the canonical surjection, then we define $\delta$ on it to be the class of $1/n$ in $\mathbb{Q}/\mathbb{Z}$. Each homomorphism $\iota:\widehat{\mu}\to\mu_{n}$ is a composite of the canonical homomorphism $\widehat{\mu}\to\mu_{n}$ and a homomorphism $\mu_{n}\to\mu_{n}$ given by $x\mapsto x^{k}$. Then we set $\delta(\iota)=k/n$. The image of $\operatorname{Hom}(\widehat{\mu},\mu_{n})$ is the group $(\frac{1}{n}\mathbb{Z}/\mathbb{Z})$.
@@
Suppose that $\mathcal{Z}$ is a closed substack of a weighted projective stack. We define a function
$|\mathcal{J}_{0}\mathcal{Z}|\to\mathbb{Q}/\mathbb{Z},\hskip 28.45274pt(x,\phi:\widehat{\mu}\to\operatorname{Aut}(f)=\mu_{n(x)})\mapsto\delta{(\phi)}.$
The function descends to a function $\gamma:\pi_{0}\mathcal{J}_{0}\mathcal{Z}\to\mathbb{Q}/\mathbb{Z}.$
@@
Suppose that $S\to\mathcal{J}_{0}\mathcal{Z}$ is a morphism with $S$ connected $F$-scheme corresponding to $(f:S\to\mathcal{Z},\phi:\widehat{\mu}\to\operatorname{Aut}(S)=\mu_{n(f)}).$ Then the value of the function at any geometric point of $S$ is $\delta(\phi)$. The claim follows. ∎
@@
Denote by $\omega:|\mathcal{J}_{0}\mathcal{Z}|\to|\mathcal{Z}|$ the canonical morphism. Given $z\in|\mathcal{Z}|$, we denote by
$(\pi_{0}\mathcal{J}_{0}\mathcal{Z})_{z}:=\{\mathcal{Y}\in\pi_{0}\mathcal{J}_{0}\mathcal{Z}|\hskip 2.84544pt\exists y\in\mathcal{Y}:\omega(y)=z\}.$
@@
Suppose $z$ is realized as $\widetilde{z}:\operatorname{Spec}(K)\to\mathcal{Z}$ with $K$ an algebraically closed $F$-field. We verify that the map is surjective: if $\mathcal{Y}\in(\pi_{0}\mathcal{J}_{0}\mathcal{Z})_{z}$, let $y\in\mathcal{Y}$ be such that $\omega(y)=z$. By increasing $K$ if needed, we realize $y$ as a $K$-point. The point $y$ then corresponds to a pair $(y_{0},f)$ of a $K$-point of $\mathcal{Z}$, necessarily isomorphic to $\widetilde{z}$, and a homomorphism $f:\widehat{\mu}\to\operatorname{Aut}(y_{0})\cong\operatorname{Aut}(\widetilde{z})=\operatorname{Aut}(z)=\mu_{\#\operatorname{Aut}(z)}$. Now it is $\delta(f)$ that maps to $\mathcal{Y}$. Let us show the injectivity. Suppose $\frac{k_{1}}{\#\operatorname{Aut}(z)}$ and $\frac{k_{2}}{\#\operatorname{Aut}(z)}$ map to the same component. The value of the function $\gamma$ at these components is $\frac{k_{1}}{\#\operatorname{Aut}(z)}$ and $\frac{k_{2}}{\#\operatorname{Aut}(z)}$, respectively. Hence $k_{1}=k_{2}$ and the function is injective. The proof is completed. ∎
@@ 3. Modular curves
$\mathscr{X}_{0}(1)\cong\mathcal{P}(4,6)$
and consequently that $\mathscr{X}_{0}(1)^{\operatorname{rig}}\cong\mathcal{P}(2,3).$ The coarse moduli space of $\mathscr{X}_{0}(1)$ is $\mathbb{P}^{1}$. The pullback of the line bundle $\mathcal{O}_{\mathbb{P}^{1}}(1)$ is the line bundle $\mathcal{O}_{\mathcal{P}(4,6)}(12)$ on $\mathscr{X}_{0}(1)$, and the Hodge bundle of $\mathscr{X}_{0}(1)$ corresponds to $\mathcal{O}_{\mathcal{P}(4,6)}(1)$, which is polarizing line bundle.
@@
Let $\mathscr{Y}_{0}(N)$ be the moduli $F$-stack of elliptic curves equipped with cyclic order $N$ subgroup schemes of $E[N]$.
@@
Clearly, one can identify $\mathscr{Y}_{0}(N)$ with the preimage of $\mathscr{Y}_{0}(1)$ in $\mathscr{X}_{0}(N)$. We denote by $Y_{0}(N)$ and by $X_{0}(N)$ the coarse moduli spaces of $\mathscr{Y}_{0}(N)$ and by $\mathscr{X}_{0}(N)$, respectively. The curves $Y_{0}(N)$ and $X_{0}(N)$ are well-known to be smooth [DR73, IV.6.7] and irreducible (see discussion in [Yan06, Section 1.1]). Thus $\mathscr{Y}_{0}(N)$ and $\mathscr{X}_{0}(N)$ are irreducible as well.
@@
Let $\Lambda$ be the Hodge line bundle on $\mathscr{X}_{0}(1)$. There exists a unique line bundle $\Lambda^{\operatorname{rig}}$ on $\mathscr{X}_{0}(1)^{\operatorname{rig}}$, such that its pullback for the rigidification morphism $\mathscr{X}_{0}(1)\to\mathscr{X}_{0}(1)^{\operatorname{rig}}$ is $\Lambda^{2}$. One has a canonical isomorphism $\mathscr{X}_{0}(1)\xrightarrow{\sim}\sqrt[2]{\Lambda^{\operatorname{rig}}/\mathscr{X}_{0}(1)^{\operatorname{rig}}}$ from Lemma 2.5 induced by $\Lambda^{2}\xrightarrow{=}(\mathscr{X}_{0}(1)\to\mathscr{X}_{0}(1)^{\operatorname{rig}})^{*}(\Lambda^{\operatorname{rig}}).$ Let us write $\Lambda_{N}^{\operatorname{rig}}:=(J_{N}^{\operatorname{rig}})^{*}(\Lambda^{\operatorname{rig}})$ and $\Lambda_{N}:=J_{N}^{*}(\Lambda)$. Let $\eta^{\operatorname{rig}}_{N}:\operatorname{Spec}(\overline{F}(X_{0}(N)))\to\mathscr{X}_{0}(N)^{\operatorname{rig}}$ be the generic point of the stacky curve $\mathscr{X}_{0}(N)^{\operatorname{rig}}$.
@@
One has that $\mathscr{X}_{0}(N)\langle F\rangle$ is Zariski-dense in $\mathscr{X}_{0}(N)$.
@@
There exists a non-empty open $U\subset X_{0}(N)\cong\mathbb{P}^{1}$, such that $\mathscr{X}_{0}(N)\cong U\times B\mu_{2}.$ As the fibers of $\mathscr{X}_{0}(N)\to X_{0}(N)$ above points in $U$ are isomorphic to $B\mu_{2}$, they have rational points. The proof follows. ∎
@@
The degree of $\mathscr{X}_{0}(N)\to\mathscr{X}_{0}(1)$ is equal to the number of cyclic subgroups of order $N$ of $(\mathbb{Z}/N\mathbb{Z})^{2}$. This number is
@@
The first part is clear. To count the number of cyclic subgroups of order $N$, we observe that is precisely $\phi(N)$ times the number of elements of order $N$ in $(\mathbb{Z}/N\mathbb{Z})^{2}$. There are precisely $N\cdot\prod_{p|N}\big(1-\frac{1}{p^{2}}\big)$ elements of order $N$ in this group. The wanted number is thus
$\frac{N\cdot\prod_{p|N}\big(1-\frac{1}{p}^{2}\big)}{N\cdot\prod_{p|N}\big(1-\frac{1}{p}\big)}=\kappa(N).$
By the fact that $\mathscr{X}_{0}(N)\to\mathscr{X}_{0}(1)$ is representable and by [Sil09, Theorem 10.1], the only points of $\mathscr{X}_{0}(N)$ where the automorphism group schemes of points may not be $\mu_{2}$ lie inside of the fibers of $\mathscr{X}_{0}(N)\to\mathscr{X}_{0}(1)$ where the $j$-value is $1728$ (for the automorphism group scheme subgroup scheme of $\mu_{4}$) and 0 (for the automorphism group scheme subgroup scheme of $\mu_{6}$). The following lemma is a slight restatement of [BV24, Lemma 3.5].
@@
We end the paragraph by proving the following lemma.
@@
    one has $J_{N}$ is étale at $\widehat{E}$ and $\operatorname{Aut}(\widehat{E})=\mu_{6}$ (respectively $\operatorname{Aut}(\widehat{E})=\mu_{4}$) if $j(J_{N}(\widehat{E}))=0$ (respectively $j(J_{N}(\widehat{E}))=1728$);
@@
where the left vertical morphism is induced from $J_{N}^{\operatorname{rig}}$ and is representable. By Lemma 3.6 this means that $m_{1}=m_{2}/\gcd(m_{2},e)$ where $e$ is the ramification degree of $R\to\widehat{R}$. By [Sch74, Page 97] one has that $e\in\{1,2,3\}$. As $m_{2}\in\{2,3\}$, this means that that $\overline{J_{N}}$ is unramified at r(E^)r($\widehat${E}) if and only if $m_{1}=m_{2}$, which gives equivalence between 1) and 3). ∎
@@
Let us start by treating the case $N=1$. The choice of the Hodge line bundle $\Lambda$ induces an identification $\mathscr{X}_{0}(1)\cong\mathcal{P}(4,6)$ with $\Lambda$ corresponding to the line bundle $\mathcal{O}(1)$. We have the untwisted sector $(x,\widehat{\mu}\xrightarrow{y\mapsto 1}\operatorname{Aut}(x)).$ We also have the twisted sector $(x,\widehat{\mu}\to\mu_{12}\xrightarrow{y\mapsto y^{6}}\mu_{2})$ for any geometric point $x$ with $\operatorname{Aut}(x)=2$. Assume that $\operatorname{Aut}(x)=6,$ hence $\operatorname{Aut}(x)=\mu_{6}$. We have $6$ such sectors by Lemma 2.15, namely the ones containing the points $([0,1],\widehat{\mu}\to\mu_{12}\xrightarrow{x\mapsto x^{2\ell}}\mu_{6})$ for $\ell=0,1,2,3,4,5$. Moreover when $\ell=0,3$, the corresponding sectors are precisely 0 and $1/2$. One treats analogously the sectors corresponding to $\operatorname{Aut}(x)=4.$ Let us calculate the ages. The anticanonical line bundle on $\mathcal{P}(4,6)$ is given by $\mathcal{O}(1)^{\otimes 10}$. The line bundle $\mathcal{O}(1)$ corresponds to the linearization $(x,y,a)\mapsto(t^{4}x,t^{6}y,t^{1}a)$ of the trivial line bundle $(\mathbb{A}^{2}-\{0\})\times\mathbb{A}^{1}\to(\mathbb{A}^{2}-\{0\})$. We deduce that the action of $\mu_{6}$, which is canonically seen as the automorphism group scheme of the point $[0,1]$, on the fiber $\mathcal{O}(1)([0,1])$ is given by $\zeta_{6}\cdot a=\zeta_{6}a.$ In other words, the action corresponds to the canonical inclusion $\mu_{6}\to\overline{F}^{*}=\mathbb{G}_{m}(\overline{F})$. Now we can calculate the age of the sector containing $([0,1],\widehat{\mu}\to\mu_{12}\xrightarrow{y\mapsto y^{2\ell}}\mu_{6})$. When $\ell=0$, the ages with respect to $\mathcal{O}(1)$ and $\mathcal{O}(1)^{\otimes 10}$ are 0. When $\gcd(\ell,6)=1$, we obtain that the induced inclusion of $\mu_{k}$ for some $k\geq 1$ to the automorphism group scheme of $[0,1]$ is given by $\mu_{6}\to\mu_{6},x\mapsto x^{\ell}$. The age with respect to $\mathcal{O}(1)$ is thus $\ell/6$. It is immediate that the age with respect to $\mathcal{O}(10)$ is $\{10\ell/6\}=\{4\ell/6\}=\{2\ell/3\}$. Assume that $\ell=2$. Then $\widehat{\mu}\to\mu_{6}$ is given by $x\mapsto x^{4}$. The induced inclusion $\mu_{k}\to\mu_{6}$ is given by $\mu_{3}\to\mu_{6},x\mapsto x$. The age with respect to $\mathcal{O}(1)$ is $1/3$, while the age with respect to $\mathcal{O}(10)=\{10/3\}=1/3$. Similarly, in the case $\ell=4$, we get that the induced morphism is $\mu_{3}\to\mu_{6},x\mapsto x^{2}$. The age with respect to $\mathcal{O}(1)$ is $2/3$, while the age with respect to $\mathcal{O}(10)$ is $2/3$. Finally, in the case $\ell=3$, we have that the induced morphism is given by $\mu_{2}\to\mu_{6}$, which has the age $1/2$ with respect to $\mathcal{O}(1)$ and the age 0 with respect to $\mathcal{O}(10)$. The case when the automorphism group is of order $4$ is treated similarly. The proof in the case $N=1$ is completed.

### v2

@@ 2.3. Sectors of some Deligne–Mumford curves
Consider a vector bundle $V$ on $X$. Given an algebraically closed $\overline{F}$-field $K$ an $x\in(\mathcal{J}_{0}X)(K)$ corresponding to $(x^{\prime},\mu_{n}\hookrightarrow\operatorname{Aut}(x^{\prime}))$, we denote by $\rho(x)$ the induced action of $\mu_{n}$ on $V(x)$. We define
@@
If $V=T_{X}$, we write $\operatorname{age}$ for $\operatorname{age}_{T_{X}}$.
@@
Let $k$ be an algebraically closed field over $\overline{F}$. A finite dimensional $k$-representation of $\widehat{\mu}$ is an $k$-homomorphism $\rho:\widehat{\mu}\to\operatorname{GL}_{r}$, for certain $r\geq 1$ called the dimension of representation, which factorizes through the canonical homomorphism $\widehat{\mu}\to\mu_{\ell}$ for some $\ell\geq 1$. Let $\rho_{\ell}:\mu_{\ell}\to\operatorname{GL}_{r}$ be the induced representation. We define
@@
Let $k$ be an algebraically closed field over $\overline{F}$ such that there exists a $k$-point $x\in\mathcal{X}$. The point $x$ corresponds to a point $x\in X(k)$ and a $k$-homomorphism $\widehat{\mu}\to\operatorname{Aut}(x).$ We have a commutative diagram
$\operatorname{Aut}(x)$$\operatorname{GL}(f^{*}V(x))=\operatorname{GL}(V(f(x)))$$\operatorname{Aut}(f(x))$$\operatorname{GL}(V(f(x)))$$\scriptstyle{=}$
@@
Suppose we have $f:X\to Y$ as in Lemma 2.12 which is also étale at a point $x$ of $X$. Let $\mathcal{X}\in\pi_{0}\mathcal{J}_{0}X$ be any sector above $x\in|X|$, and let $\mathcal{Y}=(\pi_{0}\mathcal{J}_{0}f)(\mathcal{X})$. Then
@@
As $f$ is étale at $x$, then $f^{*}T_{Y}\cong T_{X}$ on a neighborhood of $x$. Then the assertion follows from Lemma 2.12. ∎
@@ 2.3.3. Weighted projective stacks
Let $n\geq 1$ be an integer. Suppose that $a_{1},\dots,a_{n}\geq 1$ are given integers. Recall the weighted projective stack $\mathcal{P}(a_{1},\dots,a_{n}):=[(\mathbb{A}^{n}-\{0\})/\mathbb{G}_{m}]$. For any $F$-scheme $S$ and any morphism $f:S\to\mathcal{P}(a_{1},\dots,a_{n})$, we have a canonical identification $\operatorname{Aut}(f)=\mu_{n(f)}$ where $n(f)|\operatorname{lcm}(a_{i})$. Fix a bijection
@@
Observe that when $n|k$, then the canonical surjection $\widehat{\mu}\to\mu_{n}$ factors through another canonical surjection $\widehat{\mu}\to\mu_{k}$. If $\theta_{n}\colon\widehat{\mu}\to\mu_{n}$ is the canonical surjection, then let $\delta(\theta_{n})\mathrel{\mathchoice{\vbox{\hbox{$\displaystyle:$}}}{\vbox{\hbox{$\textstyle:$}}}{\vbox{\hbox{$\scriptstyle:$}}}{\vbox{\hbox{$\scriptscriptstyle:$}}}{=}}\frac{1}{n}\in\mathbb{Q}/\mathbb{Z}$. In general, a homomorphism $\iota:\widehat{\mu}\to\mu_{n}$ is a composite of $\theta_{n}$ and a homomorphism $\mu_{n}\to\mu_{n}$ given by $x\mapsto x^{k}$. Then we define $\delta(\iota)=k/n$. The image of $\operatorname{Hom}(\widehat{\mu},\mu_{n})$ via $\delta$ is the group $(\frac{1}{n}\mathbb{Z}/\mathbb{Z})$.
@@
Suppose that $\mathcal{Z}$ is a closed substack of a weighted projective stack. Consider a function
$|\mathcal{J}_{0}\mathcal{Z}|\to\mathbb{Q}/\mathbb{Z},\hskip 28.45274pt(x,\phi:\widehat{\mu}\to\operatorname{Aut}(x)=\mu_{n(x)})\mapsto\delta{(\phi)}.$
This function descends to a function $\gamma:\pi_{0}\mathcal{J}_{0}\mathcal{Z}\to\mathbb{Q}/\mathbb{Z}.$
@@
Suppose that $S\to\mathcal{J}_{0}\mathcal{Z}$ is a morphism with $S$ connected $F$-scheme corresponding to
$(f:S\to\mathcal{Z},\phi:\widehat{\mu}\to\operatorname{Aut}(f)=\mu_{n(f)}).$
Then the value of the function at any geometric point of $S$ is $\delta(\phi)$, so the claim follows. ∎
@@
Denote by $\omega:|\mathcal{J}_{0}\mathcal{Z}|\to|\mathcal{Z}|$ the canonical morphism. Given $z\in|\mathcal{Z}|$, let
$(\pi_{0}\mathcal{J}_{0}\mathcal{Z})_{z}:=\{\mathcal{Y}\in\pi_{0}\mathcal{J}_{0}\mathcal{Z}\;|\;\hskip 2.84544pt\exists y\in|\mathcal{Y}|\text{ s.t. }\omega(y)=z\}.$
@@
For surjectivity of $\gamma$, suppose that $z$ is realized as $\widetilde{z}:\operatorname{Spec}(K)\to\mathcal{Z}$ with $K$ an algebraically closed $F$-field. We verify that the map is surjective: if $\mathcal{Y}\in(\pi_{0}\mathcal{J}_{0}\mathcal{Z})_{z}$, let $y\in\mathcal{Y}$ be such that $\omega(y)=z$. By increasing $K$ if needed, we realize $y$ as a $K$-point. The point $y$ then corresponds to a pair $(y_{0},f)$ of a $K$-point of $\mathcal{J}_{0}\mathcal{Z}$, where $y_{0}\cong\widetilde{z}$ and $f:\widehat{\mu}\to\operatorname{Aut}(y_{0})\cong\operatorname{Aut}(\widetilde{z})=\operatorname{Aut}(z)=\mu_{\#\operatorname{Aut}(z)}$. Hence $\delta(f)$ maps to $\mathcal{Y}$, proving surjectivity.
For injectivity, suppose that $\frac{k_{1}}{\#\operatorname{Aut}(z)}$ and $\frac{k_{2}}{\#\operatorname{Aut}(z)}$ map to the same component. The value of $\gamma$ at these components is $\frac{k_{1}}{\#\operatorname{Aut}(z)}$ and $\frac{k_{2}}{\#\operatorname{Aut}(z)}$, respectively. Hence $k_{1}=k_{2}$ and so $\gamma$ is injective. ∎
@@ 3. Modular curves
$\mathscr{X}_{0}(1)\cong\mathcal{P}(4,6),$
so $\mathscr{X}_{0}(1)^{\operatorname{rig}}\cong\mathcal{P}(2,3)$. Its coarse moduli space is $X_{0}(1)\cong\mathbb{P}^{1}$. The pullback of the line bundle $\mathcal{O}_{\mathbb{P}^{1}}(1)$ is the line bundle $\mathcal{O}_{\mathcal{P}(4,6)}(12)$ on $\mathscr{X}_{0}(1)$, and the Hodge bundle $\Lambda$ on $\mathscr{X}_{0}(1)$ corresponds to $\mathcal{O}_{\mathcal{P}(4,6)}(1)$, which is a polarizing line bundle.
@@
Let $\mathscr{Y}_{0}(N)$ be the moduli $F$-stack of elliptic curves equipped with cyclic order $N$ subgroup schemes.
@@
Clearly, one can identify $\mathscr{Y}_{0}(N)$ with the preimage of $\mathscr{Y}_{0}(1)$ in $\mathscr{X}_{0}(N)$. We denote by $Y_{0}(N)$ and $X_{0}(N)$ the coarse moduli spaces of $\mathscr{Y}_{0}(N)$ and $\mathscr{X}_{0}(N)$ respectively. The stacks $\mathscr{Y}_{0}(N)$ and $\mathscr{X}_{0}(N)$ are smooth by [DR73, IV.6.7], and their coarse spaces are irreducible (see discussion in [Yan06, Section 1.1]). Thus $\mathscr{Y}_{0}(N)$ and $\mathscr{X}_{0}(N)$ are irreducible as well.
@@
There exists a unique line bundle $\Lambda^{\operatorname{rig}}$ on $\mathscr{X}_{0}(1)^{\operatorname{rig}}$, such that its pullback for the rigidification morphism $\mathscr{X}_{0}(1)\to\mathscr{X}_{0}(1)^{\operatorname{rig}}$ is $\Lambda^{2}$. One has a canonical isomorphism $\mathscr{X}_{0}(1)\xrightarrow{\sim}\sqrt[2]{\Lambda^{\operatorname{rig}}/\mathscr{X}_{0}(1)^{\operatorname{rig}}}$ from Lemma 2.5 induced by $\Lambda^{2}\xrightarrow{=}(\mathscr{X}_{0}(1)\to\mathscr{X}_{0}(1)^{\operatorname{rig}})^{*}(\Lambda^{\operatorname{rig}}).$ Let $\Lambda_{N}^{\operatorname{rig}}:=(J_{N}^{\operatorname{rig}})^{*}(\Lambda^{\operatorname{rig}})$ and $\Lambda_{N}:=J_{N}^{*}(\Lambda)$. Let $\eta^{\operatorname{rig}}_{N}:\operatorname{Spec}(\overline{F}(X_{0}(N)))\to\mathscr{X}_{0}(N)^{\operatorname{rig}}$ be the generic point of the stacky curve $\mathscr{X}_{0}(N)^{\operatorname{rig}}$.
@@
The set $\mathscr{X}_{0}(N)\langle F\rangle$ is Zariski-dense in $\mathscr{X}_{0}(N)$.
@@
There exists a non-empty open $U\subset X_{0}(N)\cong\mathbb{P}^{1}$, such that $\mathscr{X}_{0}(N)\times_{X_{0}(N)}U\cong U\times B\mu_{2}$. As $U\langle F\rangle$ is Zarski-dense in $U$, the assertion follows. ∎
@@
The degree of $J_{N}\colon\mathscr{X}_{0}(N)\to\mathscr{X}_{0}(1)$ is equal to the number of cyclic subgroups of order $N$ of $(\mathbb{Z}/N\mathbb{Z})^{2}$. This number is
@@
The first part is clear. Also, the number of cyclic subgroups of order $N$ is equal to the index $\operatorname{SL}_{2}(\mathbb{Z}):\Gamma_{0}(N)]$ where $\Gamma_{0}(N)$ consists of elements of $\operatorname{SL}_{2}(\mathbb{Z})$ that’s upper-triangular in $\operatorname{SL}_{2}(\mathbb{Z}/N\mathbb{Z})$. This number is exactly $\kappa(N)$ as above by [Shi71, Proposition 1.43]. ∎
Because $\mathscr{X}_{0}(N)\to\mathscr{X}_{0}(1)$ is representable, by [Sil09, Theorem 10.1], the only points of $\mathscr{X}_{0}(N)$ where the stabilizer may not be $\mu_{2}$ lie inside the fibers of $\mathscr{X}_{0}(N)\to\mathscr{X}_{0}(1)$ where the $j$-invariant is 0 or $1728$; in those cases, the stabilizer group can be $\mu_{6}$ or $\mu_{4}$ respectively. The following lemma is a slight restatement of [BV24, Lemma 3.5], where $\sqrt[n]{\operatorname{Spec}R}$ for a DVR $R$ with the special point 0 means $\sqrt[n]{0/\operatorname{Spec}R}$.
@@
We end the paragraph by proving the following lemmas.
@@
    $J_{N}$ is étale at $\widehat{E}$ and $\operatorname{Aut}(\widehat{E})=\mu_{6}$ (respectively $\operatorname{Aut}(\widehat{E})=\mu_{4}$) if $j(J_{N}(\widehat{E}))=0$ (respectively $j(J_{N}(\widehat{E}))=1728$);
@@
where the left vertical morphism is induced from $J_{N}^{\operatorname{rig}}$ and so representable. By Lemma 3.6, this means that $m_{1}=m_{2}/\gcd(m_{2},e)$ where $e$ is the ramification degree of $R\to\widehat{R}$. By [Sch74, Page 97], one has that $e\in\{1,2,3\}$. As $m_{2}\in\{2,3\}$, this means that that $\overline{J_{N}}$ is unramified at r(E^)r($\widehat${E}) if and only if $m_{1}=m_{2}$, which gives equivalence between (1) and (3). ∎
@@
First, consider $N=1$ case. Recall that the Hodge bundle $\Lambda$ induces an identification $\mathscr{X}_{0}(1)\cong\mathcal{P}(4,6)$ with $\Lambda$ corresponding to $\mathcal{O}(1)$. The untwisted sector is $(x,\widehat{\mu}\xrightarrow{y\mapsto 1}\operatorname{Aut}(x))$. We also have the twisted sector $(x,\widehat{\mu}\to\mu_{12}\xrightarrow{y\mapsto y^{6}}\mu_{2})$ for any geometric point $x$ with $\operatorname{Aut}(x)=2$. Assume that $\operatorname{Aut}(x)=6,$ hence $x=[0,1]$ and $\operatorname{Aut}(x)=\mu_{6}$. We have $6$ such sectors by Lemma 2.15, namely the ones containing $(x,\widehat{\mu}\to\mu_{12}\xrightarrow{x\mapsto x^{2\ell}}\mu_{6})$ for $\ell=0,1,2,3,4,5$. When $\ell=0,3$, $\gamma$-values of the corresponding sectors are precisely 0 and $1/2$. One treats analogously the sectors corresponding to $\operatorname{Aut}(x)=4$.
Let’s calculate the ages. The anticanonical line bundle on $\mathcal{P}(4,6)$ is $\mathcal{O}(1)^{\otimes 10}$. The linearization describing $\mathcal{O}(1)$ is $(x,y,a)\mapsto(t^{4}x,t^{6}y,t^{1}a)$ on the trivial line bundle $(\mathbb{A}^{2}-\{0\})\times\mathbb{A}^{1}\to(\mathbb{A}^{2}-\{0\})$. Then the action of the stabilizer $\mu_{6}$ of $[0,1]$ on the fiber $\mathcal{O}(1)([0,1])$ is given by $\zeta_{6}\cdot a=\zeta_{6}a$; this action corresponds to the canonical inclusion $\mu_{6}\to\overline{F}^{*}=\mathbb{G}_{m}(\overline{F})$. Now let’s calculate the age of the sector containing $([0,1],\widehat{\mu}\to\mu_{12}\xrightarrow{y\mapsto y^{2\ell}}\mu_{6})$. When $\ell=0$, the ages with respect to $\mathcal{O}(1)$ and $\mathcal{O}(1)^{\otimes 10}$ are 0. When $\gcd(\ell,6)=1$, then the induced inclusion of $\mu_{k}$ for some $k\geq 1$ to the stabilizer of $[0,1]$ is given by $\mu_{6}\to\mu_{6},x\mapsto x^{\ell}$. The age with respect to $\mathcal{O}(1)$ is thus $\ell/6$. It is immediate that the age with respect to $\mathcal{O}(10)$ is $\{10\ell/6\}=\{4\ell/6\}=\{2\ell/3\}$. Assume that $\ell=2$. Then $\widehat{\mu}\to\mu_{6}$ is given by $x\mapsto x^{2}$. The induced inclusion $\mu_{k}\to\mu_{6}$ is given by $\mu_{3}\to\mu_{6},x\mapsto x$. The age with respect to $\mathcal{O}(1)$ is $1/3$, while the age with respect to $\mathcal{O}(10)=\{10/3\}=1/3$. Similarly, when $\ell=4$, the induced morphism is $\mu_{3}\to\mu_{6},x\mapsto x^{2}$. The age with respect to $\mathcal{O}(1)$ is $2/3$, while the age with respect to $\mathcal{O}(10)$ is $2/3$. Finally, when $\ell=3$, the induced morphism is $\mu_{2}\to\mu_{6}$, which has the age $1/2$ with respect to $\mathcal{O}(1)$ and the age 0 with respect to $\mathcal{O}(10)$. The case when the automorphism group is of order $4$ is treated similarly. The proof in the case $N=1$ is completed.